40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Mar 29 18:27:27 2017, awade, DailyProgress, BEAT, Reconstructing free running laser beat note detector 
    Reply  Thu Mar 30 15:49:51 2017, awade, DailyProgress, BEAT, Reconstructing free running laser beat note detector 
Message ID: 1831     Entry time: Wed Mar 29 18:27:27 2017     Reply to this: 1832
Author: awade 
Type: DailyProgress 
Category: BEAT 
Subject: Reconstructing free running laser beat note detector 

We now have >150 kHz UGF on both loops. This should be enough for an initial beat note measurement. The improved UGF in the south path was because I switch the SRS DB64 delay line box in the PD RF electronic signal path to the FSS box for a fixed cable length.  I didn't really look further into why the delayer box was causing this, once it was solved I moved on. Maybe internal reflections or loss. No bode plots/spectra/traces here, we'll do these properly again when Craig is finished with his candidacy presentation.  

On Monday we locked both cavities with pre-cavity beat note tuned to within the 125 MHz bandwidth of the NF1811 (125 MHz, InGaAs) transmission BN detector. We saw nothing for transmitted detector.  

In the pre-cavity BN dector we installed the new Newfocus 1611 detector (the other was returned to the 40m) and Craig and I were able to just find a faint beat note on the pre-cavity BN detector.  It was about 5 dB above the noise floor of the spectrum analyzer, so not great.

Beat note show up (~160 MHz) with south laser slow voltage ~0.5 - 6 V and north laser slow voltage of 3.0 V.  

Originally I'd spend no more than two hours setting this up and finding a beat note to start with so beam lines were not squared with the table and it generally needed neating up: it needed a rebuild anyway.  I spend some of today rebuilding.  Even though the overlap of the beams looked good and there was ~0.5 mW of light from each laser I was unable to find the beat note again. Its not very well mode matched but we would still expect at least some overlap.  It could also be that I need to step more slowly again, its always initially hard to get a fix on it.  Will try again tomorrow.

 

Ordering:

Also, we some how ordered the metric mounted version of the Newfocus 1611, we only realized this Monday.  I've emailed them to find out if they can just sell us the new shoe, they suggested ordering a converter grub screw TA-8M4-10, might be the cheapest option.

ELOG V3.1.3-